OpenStudy (anonymous):

solve the system by graphing -3x-y=-10 4x-4y=8

OpenStudy (anonymous):

-3x-y=-10 plug 0 for x -y=-10 --> y=10 so we get (0,10) plug 1 for x -3-y=-10 -y=-7 y=7 (1,7) we have two points (0,10) and (1,7) connects the dots